# Theoretical Bounds for Stable In-Context Learning

### v1

@@ 3.5 A two-stage observable (plug-in) scheme and finite-sample guarantee
1:tolerance $\xi\in(0,1)$, target $\delta\in(0,1)$, initial $K_{0}$ (e.g., $K_{0}=100$)
@@
3:  Draw $K_{0}$ demos, compute $\widehat{\Sigma}_{K_{0}}$, set $\widehat{\lambda}_{0}=\lambda_{\min}(\widehat{\Sigma}_{K_{0}})$, $\widehat{\|\Sigma\|}=\|\widehat{\Sigma}_{K_{0}}\|$, $\widehat{r}_{\mathrm{eff}}=\mathrm{tr}(\widehat{\Sigma}_{K_{0}})/\|\widehat{\Sigma}_{K_{0}}\|$.
4:  Lower confidence bound (use (9) constant $C$):
@@
$\underline{\lambda}=\widehat{\lambda}_{0}-C\,\widehat{\|\Sigma\|}\!\left(\sqrt{\tfrac{\widehat{r}_{\mathrm{eff}}\log(4/\xi)}{K_{0}}}+\tfrac{\widehat{r}_{\mathrm{eff}}\log(4/\xi)}{K_{0}}\right).$
5:  if $\underline{\lambda}\leq 0$ then
6:   $K_{0}\leftarrow 2K_{0}$ and repeat
7:  end if
8:until $\underline{\lambda}>0$
9:Set $\widehat{\Delta}=\underline{\lambda}-\delta$.
10:Compute the second-stage size using the plug-in form of (10):
$K\;\;=\;\;C^{\prime}\,\frac{\widehat{\|\Sigma\|}^{2}\,\widehat{r}_{\mathrm{eff}}\log(2/\xi)}{\widehat{\Delta}^{2}}.$
11:return total $K_{\mathrm{final}}=K_{0}+K$.
3.6 Final implementable theorem statement
[ICL Minimal Necessary Length, non-asymptotic, observable two-stage implementation] Under the above setup, let $\delta\in(0,\lambda_{\min}(\Sigma))$ and failure tolerance $\xi\in(0,1)$. There exist absolute constants $C_{1},C_{2},C_{3}$ (depending on matrix-Bernstein and sub-Gaussian constants) such that:
  1. 1.
    (Non-observable bound using spectral quantities) If $\Delta=\lambda_{\min}(\Sigma)-\delta>0$ and
    then $\Pr(\lambda_{\min}(\widehat{\Sigma}_{K})\geq\delta)\geq 1-\xi$. Ignoring lower-order terms (and using the sharper operator-norm concentration) yields the simplified dominant form
    $K\geq C_{2}\frac{\|\Sigma\|^{2}r_{\mathrm{eff}}\log(2/\xi)}{(\lambda_{\min}(\Sigma)-\delta)^{2}},$
    which suffices in practice.
  2. 2.
    then forming $\widehat{\Sigma}_{K_{0}}$ and constructing the lower confidence bound $\lambda$ as in (11), and then setting $\widehat{\Delta}=\lambda-\delta$ and choosing $K$ according to the rule in item (1) (with plug-in estimates $\|\Sigma\|_{\wedge},\ \widehat{r}_{\mathrm{eff}}$) yields, with total failure probability $\leq\xi$, the guarantee $\lambda_{\min}(\widehat{\Sigma}_{K_{0}+K})\geq\delta$.
3.7 Connection to ICL (in-context learning) and experimental suggestions
A one-shot practical calibration.
To tighten $K^{*}$ for practice, we (i) replace $\|\Sigma\|$ by $\|\Sigma\|_{\mathrm{F}}$, (ii) use the trace-based effective rank $r_{\mathrm{eff}}^{\mathrm{tr}}=\mathrm{tr}(\Sigma)^{2}/\|\Sigma\|_{\mathrm{F}}^{2}$, and (iii) substitute the minimum eigenvalue by a $q$-quantile eigenvalue $\lambda_{q}$ ($q\in[0.05,0.2]$ tuned on a small validation set; no test leakage). We then compute
$K^{\mathrm{cal}}\;=\;\alpha\cdot\frac{\|\widehat{\Sigma}\|_{\mathrm{F}}^{\,2}\;r_{\mathrm{eff}}^{\mathrm{tr}}\;\log(2/\xi)}{(\lambda_{q}-\delta)^{2}},$
where a single global scale $\alpha$ is estimated once per encoder–generator pair from a small validation subset (e.g., $5$ tasks). This variant preserves the same monotone dependence as the theory while substantially reducing over-conservativeness.
Theoretical meaning
Using $\lambda_{\min}(\widehat{\Sigma}_{K})$ as a proxy for “ICL learnability/stability” is reasonable because the condition number (and hence predictive variance in ridge regression or linear approximation) depends on inverse eigenvalues. For example, prediction variance scales with $\phi(x^{\ast})^{\top}(H_{K}^{\top}H_{K}+\lambda I)^{-1}\phi(x^{\ast})$, which is proportional to $(K\lambda_{\min}(\widehat{\Sigma}_{K})+\lambda)^{-1}$. If $\lambda_{\min}(\widehat{\Sigma}_{K})$ is lower bounded by $\delta$, then prediction variance is upper bounded and ICL is more likely to behave stably.
Empirical validation of assumptions.
Although our main theorem is derived under the nominal assumption (A1), we conduct diagnostic experiments on hidden representations from GPT-2 models to evaluate how well these assumptions hold in practice. We find that (i) hidden states are approximately sub-Gaussian at intermediate layers, though heavier tails appear in later layers; (ii) representation drift across prompt positions is present but moderate; (iii) the spectral statistic $\lambda_{\min}(\hat{\Sigma}_{K})$ correlates with empirical ICL stability. These results indicate that our theoretical assumptions serve as a reasonable proxy, and the relaxed extensions in Appendix D capture the observed deviations.
Synthetic stress tests.
To further validate our theoretical claims, we perform controlled synthetic stress tests (Appendix E) by introducing bounded drift, heavy-tailed features, and temporal dependence. Across all cases, we find that empirical knee-points increase smoothly with the violation strength, in line with our robustness extensions, confirming that the lower bound remains a conservative but reliable predictor of ICL stability.
Remark on interpretation.
While we use $\lambda_{\min}(\hat{\Sigma}_{K})$ as a stability proxy, we recognize that it can be conservative and may not fully capture all sources of instability (e.g. attention saturation). Appendix F provides robustness analyses, alternative spectral surrogates, and empirical diagnostics that validate and calibrate this proxy.
4 Experiments
We conduct experiments to validate the theoretical lower bound on the minimal prompt length required for stable In-Context Learning (ICL). The experiments test the predictions of our theory across various datasets, encoders, and generators.
4.1 Measuring the Knee-Point
The knee-point $\widehat{K}_{\text{knee}}$ is defined as the point of minimal sum of squared errors (SSE) in a two-segment piecewise linear fit of the accuracy curve $\{(K,\mathrm{ACC}(K))\}_{K\in\mathcal{K}}$. We calculate the knee-point by selecting the split index $t$ that minimizes SSE, and the uncertainty is quantified via bootstrapping with 200 resamplings, reporting the 95% percentile interval for $\widehat{K}_{\text{knee}}$.
Table 1: Error ratio ($K^{*}/\text{knee}$) across datasets, encoders, and generators. Each number is averaged over 5 random seeds. Lower is better (closer to 1). The Raw column uses the default parameter values for prediction, whereas the Calibrated column reflects a few exploratory adjustments to some of these values; the procedure for determining each parameter is detailed in H.2.
Dataset Encoder GPT-2 small (124M) GPT-2 medium (345M) GPT-3 (175B)
Raw Calibrated Raw Calibrated Raw Calibrated
@@ 4.1 Measuring the Knee-Point
4.2 Results Across Datasets
Table 1 summarizes the empirical knee-points across datasets. Theoretical predictions consistently exceed empirical knee-points with an average error ratio of 2.9×, confirming the reliability of the theoretical upper bound. The calibrated estimates ($K_{\mathrm{cal}}$) track the knee-point more closely.
4.3 Task-Dependence of Error Ratios
The error ratio varies across datasets. More complex tasks, such as AGNews, require larger prompts compared to simpler tasks like TREC-6. This aligns with our theory, which predicts larger prompts for more complex tasks.
4.4 Model Scale Effect
We evaluate the effect of model scale on ICL performance by testing a 124M and a 345M model. Larger models reduce the empirical knee-point, and the error ratio improves. For SST-2, the knee decreases from $\approx$12 for the 124M model to $\approx$7 for the 345M model, with the error ratio decreasing from 3.4× to 2.2×.
4.5 Impact of Tokenization and Prompt Truncation
The choice of tokenization and prompt truncation policies significantly affects ICL stability. Single-token labels (e.g., "yes/no") lead to smaller knees and error ratios (1.2–1.5×), whereas multi-token labels (e.g., "positive/negative") result in error ratios greater than 3×. For truncation, "latest-first" truncation produces the smallest knees and error ratios (1.2–1.8×), while "earliest-first" and "random" truncations lead to larger error ratios.
4.6 Synthetic Stress Tests
We further evaluate the robustness of our theoretical bound by introducing controlled violations of the nominal assumption (A1). These stress tests involve the following: Bounded Drift: As the drift parameter $\epsilon$ increases, the empirical knee-point increases linearly. This confirms the additive correction predicted by our robustness analysis in B. Heavy-Tailed Features: As the tail parameter $\nu$ decreases, the empirical knee-point increases multiplicatively. However, truncation or normalization significantly reduces this inflation, as detailed in D. Ablation Studiesis shown in G.
These observations confirm that the lower bound remains a reliable predictor of ICL stability across different data violations.

### v2

@@ 3.3 An observable two-stage estimator
0: tolerance $\xi\in(0,1)$; target $\delta$; ridge coefficient $\epsilon_{\rho}>0$; initial $K_{0}$ (e.g., $K_{0}=100$)
[…]
@@
3:  $\xi_{j}\leftarrow\xi/2^{j+2}$ {geometric allocation; $\sum_{j\geq 0}\xi_{j}\leq\xi/2$}
4:  Draw $K_{0}$ demonstrations and compute $\widehat{\Sigma}_{K_{0}}$
5:  $\rho\leftarrow\min\{\epsilon_{\rho}\cdot\mathrm{tr}(\widehat{\Sigma}_{K_{0}})/d,\ \delta/2\}$
6:  $\widehat{\Sigma}_{K_{0},\rho}\leftarrow\widehat{\Sigma}_{K_{0}}+\rho I_{d}$
7:  $\widehat{\lambda}_{0}\leftarrow\lambda_{\min}(\widehat{\Sigma}_{K_{0},\rho})$
8:  $\widehat{\|\Sigma\|}\leftarrow\|\widehat{\Sigma}_{K_{0}}\|$
9:  $\widehat{r}_{\mathrm{eff}}\leftarrow\mathrm{tr}(\widehat{\Sigma}_{K_{0}})/\|\widehat{\Sigma}_{K_{0}}\|$
10:  Lower confidence bound:
@@
11:  $\underline{\lambda}\leftarrow\widehat{\lambda}_{0}-C\,\widehat{\|\Sigma\|}\!\left(\sqrt{\tfrac{\widehat{r}_{\mathrm{eff}}\log(2/\xi_{j})}{K_{0}}}+\tfrac{\widehat{r}_{\mathrm{eff}}\log(2/\xi_{j})}{K_{0}}\right)$
12:  if $\underline{\lambda}\leq 0$ then
13:   $K_{0}\leftarrow 2K_{0}$; $j\leftarrow j+1$
14:  end if
15:until $\underline{\lambda}>0$
[…]
17:if $\widehat{\Delta}\leq 0$ then
18:  return infeasible target
19:end if
[…]
21:return $K_{\mathrm{final}}\leftarrow K_{0}+K$
Algorithm 1 is fully observable given a feature map $\phi$: it requires only computing Gram/covariance statistics of demonstration representations. If it returns infeasible target, then the requested floor $\delta$ exceeds the current lower confidence bound $\underline{\lambda}$ (i.e., $\widehat{\Delta}\leq 0$), indicating that the chosen target is not supported at confidence $1-\xi$ under the current feature map and ridge level; practical remedies and calibration guidance are given in Appendix J. The appendix contains a complete non-asymptotic analysis showing how to choose $K_{0}$ and $\eta$ (and how to split $\xi$) so that the returned $\widehat{K}$ attains the target guarantee in (5) under the same assumptions as Proposition 3.3.
Both stages operate solely on the demonstration pool (or training split) and do not access test-query outputs or labels; the resulting prompt length is fixed prior to evaluation.
3.4 Spectral floor as stability proxy
The proxy in (2) is purely geometric: it asks whether the demonstration features span the space well enough so that the empirical second moment is well-conditioned. This section provides a modeling rationale—via ridge-regularized linearization—for using such a spectral “floor” as a surrogate for ICL stability, and clarifies what the proxy does and does not capture. We complement the theory with empirical diagnostics in Appendix H and direct stability measurements in Appendix L.
Many analyses of ICL posit that, for a fixed query $x_{q}$, the model’s output depends on demonstrations primarily through a low-dimensional statistic built from representations (e.g., attention-weighted averages, or a fitted linear probe). A generic abstraction is: (i) the prompt induces a parameter estimate $\widehat{w}$ by solving a regularized least-squares problem in feature space, and (ii) the model output at the query is a Lipschitz function of $\widehat{w}$. Under this view, stability reduces to controlling how much $\widehat{w}$ varies when we resample the demonstrations.
Let $H_{K}\in\mathbb{R}^{K\times d}$ be the feature matrix defined in Section 3.1. Consider the ridge estimator
$\widehat{w}(S_{K})\;:$ $=\;\arg\min_{w\in\mathbb{R}^{d}}\;\frac{1}{K}\|H_{K}w-y\|_{2}^{2}+\rho\|w\|_{2}^{2}$
$=\;\big(\widehat{\Sigma}_{K,\rho}\big)^{-1}\,\Big(\frac{1}{K}H_{K}^{\top}y\Big),$ (9)
where $y\in\mathbb{R}^{K}$ denotes supervision associated with the demonstrations (e.g., class labels under a fixed encoding, or pseudo-labels produced by the model itself), and $\rho\geq 0$ is the same ridge level as in (2).
Why spectral coverage implies stability.
At a high level, the smallest eigenvalue of the empirical second-moment matrix $\widehat{\Sigma}_{K}=\frac{1}{K}\sum_{i=1}^{K}\phi(x_{i})\phi(x_{i})^{\top}$ quantifies how well the representation vectors of the demonstrations span the feature space. When $\lambda_{\min}(\widehat{\Sigma}_{K})$ is small, the induced inverse geometry is poorly conditioned: small perturbations in the empirical moments—arising from changing or resampling a few demonstrations—can be greatly amplified, leading to large variations in the fitted coefficients and, consequently, in the model’s predictions. In contrast, a larger spectral lower bound ensures that all directions in representation space are sufficiently covered, which stabilizes the inverse problem and suppresses amplification effects. This intuition extends to quantile-based spectral surrogates, which trade off strict worst-case control for improved numerical robustness in finite samples.
Lemma 3.4 (Spectral floor $\Rightarrow$ bounded sensitivity of ridge parameters).
Assume $\|y\|_{2}\leq B\sqrt{K}$ almost surely for some $B>0$. Let $S_{K}$ and $S^{\prime}_{K}$ be two independent $K$-demo prompts with feature matrices $H_{K},H^{\prime}_{K}$ and ridge covariances $\widehat{\Sigma}_{K,\rho},\widehat{\Sigma}^{\prime}_{K,\rho}$. On the event
$\mathcal{E}_{\delta}\;:=\;\Big\{\lambda_{\min}(\widehat{\Sigma}_{K,\rho})\geq\delta\Big\}\;\cap\;\Big\{\lambda_{\min}(\widehat{\Sigma}^{\prime}_{K,\rho})\geq\delta\Big\},$
the ridge solutions satisfy the deterministic bound
$\|\widehat{w}(S_{K})-\widehat{w}(S^{\prime}_{K})\|_{2}\;\leq\;\frac{2B}{\delta}\;+\;\frac{B}{\delta^{2}}\,\|\widehat{\Sigma}_{K}-\widehat{\Sigma}^{\prime}_{K}\|\,.$ (10)
In particular, conditioning improves as the spectral floor $\delta$ increases.
From ridge sensitivity to distributional stability (sketch). Lemma 3.4 shows that a spectral floor $\lambda_{\min}(\widehat{\Sigma}_{K,\rho})\geq\delta$ prevents the ridge inverse from amplifying prompt-to-prompt variability in the induced ridge parameter $\widehat{w}$. Under a mild Lipschitz dependence of the model output distribution on $\widehat{w}$, this translates into small divergence between $p_{\theta}(\cdot\mid P_{K})$ across independently resampled prompts, with probability controlled by the proxy failure level $\xi$. We provide the full derivation (including the Lipschitz condition, intermediate inequalities, and sharper variants that avoid explicit $\|\widehat{\Sigma}_{K}-\widehat{\Sigma}^{\prime}_{K}\|$ terms) in Appendix H.
3.5 Calibration and practical recommendations
The bound in Proposition 3.3 and the plug-in rule (8) are intentionally conservative: they protect against worst-case directions in the feature space and use high-probability concentration. In practice, two issues commonly arise: (i) the smallest eigenvalue is overly sensitive to noise in near-null directions, and (ii) constants hidden in non-asymptotic inequalities can inflate the suggested prompt length. This section summarizes lightweight calibration strategies that preserve the monotonic behavior of the proxy while producing practically useful $K$.
Instead of targeting the extreme floor $\lambda_{\min}$, it is often more stable to track the quantile eigenvalue $\lambda_{q}(\widehat{\Sigma}_{K,\rho})$ defined in Section 3.1. Operationally, one may replace the event in (2) by
$\Pr\!\big(\lambda_{q}(\widehat{\Sigma}_{K,\rho})\geq\delta_{q}\big)\;\geq\;1-\xi,$ (11)
for a chosen $q\in(0,1)$ (e.g., $q\in[0.05,0.2]$) and threshold $\delta_{q}$. This reduces sensitivity to a small number of poorly-covered directions while still enforcing broad spectral coverage. The same two-stage estimator (Algorithm 1) applies by replacing $\widehat{\lambda}_{0}$ in (6) with $\lambda_{q}(\widehat{\Sigma}_{K_{0}})$ and $\Delta_{\rho}$ in (3) with the corresponding quantile margin.
Let $\widehat{K}$ denote the plug-in output from (8). We optionally apply a global calibration factor $\alpha\in(0,1]$ and use
$\widehat{K}_{\mathrm{cal}}\;:=\;\left\lceil\alpha\,\widehat{K}\right\rceil$. (12)
A practical alternative (used in our experiments) is to calibrate $(q,\alpha)$ against empirical knee-points estimated on training data only (no test leakage; Appendix C). For each encoder–generator pair, we select a small set of calibration tasks and estimate $\widehat{K}_{\mathrm{knee}}^{(t)}$ on a held-out validation subset drawn from the training split. We then choose $q^{\star}\in[0.05,0.2]$ by minimizing the median absolute log-error between $\widehat{K}_{\mathrm{raw}}^{(t)}(q)$ and $\widehat{K}_{\mathrm{knee}}^{(t)}$ across calibration tasks, and set $\alpha$ by a robust median ratio fit:
$\alpha\;:=\;\mathrm{median}_{t}\ \widehat{K}_{\mathrm{knee}}^{(t)}\big/\widehat{K}_{\mathrm{raw}}^{(t)}(q^{\star}).$
Finally, we freeze $(q^{\star},\alpha)$ and report both raw and calibrated predictions on disjoint evaluation tasks.
Default choices.
Across tasks, we recommend: (i) a modest ridge $\rho$ to stabilize inversion and dampen near-null directions; (ii) choosing $\delta$ (or $\delta_{q}$) as a fraction of the pilot spectrum, e.g., proportional to $\lambda_{q}(\widehat{\Sigma}_{K_{0},\rho})$; and (iii) splitting the failure budget evenly across the two stages (as in Algorithm 1) unless one stage is known to dominate the uncertainty. These choices are heuristic but align with the role of $\Delta_{\rho}$ in (3): larger margins reduce the required $K$, and ridge improves numerical stability when $\Sigma$ is ill-conditioned.
Table 1: Error ratio ($K^{*}/\text{knee}$) across datasets and task families for commercial API generators. Encoder groups a/b/c correspond to all-mpnet-base-v2, all-MiniLM-L6-v2, and paraphrase-MiniLM-L12-v2. Entries report the mean ratio over $5$ random seeds; per-cell standard deviations (and per-seed raw outputs) are provided in the released artifact and summarized in Appendix C. Knee-points are estimated on the test split via a two-segment piecewise-linear fit on the grid $\mathcal{K}$ with bootstrap confidence intervals (Appendix C). The Raw column uses the default theoretical constants, whereas Cal. uses validation-only calibration (Appendix J.2). Lower is better (closer to 1).
Dataset Enc.
OpenAI gpt-5.2-chat-latest
Anthropic claude-sonnet-4-5
Gemini gemini-3-flash-preview
DeepSeek deepseek-reasoner
Kimi kimi-k2-thinking
Qwen qwen-max-latest
Raw Cal. Raw Cal. Raw Cal. Raw Cal. Raw Cal. Raw Cal.
@@ 3.5 Calibration and practical recommendations
Task family Dataset Raw Cal. Raw Cal. Raw Cal. Raw Cal. Raw Cal. Raw Cal.
[…]
4 Experiments
We conduct experiments to validate the theoretical lower bound on the minimal prompt length required for stable In-Context Learning (ICL). The experiments test the predictions of our theory across various datasets, encoders, and generators. Unless otherwise stated, we use deterministic decoding (e.g., temperature =0=0 when supported) so that variability is dominated by demonstration resampling rather than token-level sampling noise. Full reproducibility specifications (prompt templates, demo selection/retrieval, representation extraction, decoding hyperparameters, random seeds, knee-point estimator, and confidence intervals) are provided in Appendix C.
4.1 Experimental setup
We provide a comprehensive experimental inventory to facilitate reproduction, ablation and future benchmarking:
Datasets: SST-2(Socher et al., 2013), AGNews(Zhang et al., 2015), Yahoo!Answers(Zhang et al., 2015), TREC-6(Voorhees and Harman, 1998), GSM8K(Cobbe et al., 2021), SVAMP(Patel et al., 2021), HotpotQA(Yang et al., 2018) and Natural Questions(Kwiatkowski et al., 2019), covering sentiment, news topic, QA and multi-step reasoning.
Encoders: all-mpnet-base-v2, all-MiniLM-L6-v2 and paraphrase-MiniLM-L12-v2(Reimers and Gurevych, 2019; Song et al., 2020; Wang et al., 2021).
Models: Anthropic Claude-Sonnet-4-5, DeepSeek-Reasoner(Guo et al., 2025), Gemini-3-Flash-Preview, Kimi-k2-thinking(Team et al., 2025), OpenAI gpt-5.2-chat-latest and Qwen-Max-Latest(Yang et al., 2025).
4.2 Measuring the Knee-Point
For each (task, encoder, generator) setting, we estimate an empirical saturation point $\widehat{K}_{\mathrm{knee}}$ from an accuracy–vs.–prompt-length curve evaluated on a discrete grid $\mathcal{K}$. Concretely, for each $K\in\mathcal{K}$ we measure $\mathrm{ACC}(K)$ on the test split under a fixed demonstration construction protocol (including retrieval, formatting, and decoding). We then fit a two-segment piecewise-linear model to $\{(K,\mathrm{ACC}(K))\}_{K\in\mathcal{K}}$ and define $\widehat{K}_{\mathrm{knee}}$ as the breakpoint minimizing the sum of squared errors (SSE), with ties broken toward smaller $K$ for conservatism. To quantify uncertainty, we compute a percentile bootstrap confidence interval for $\widehat{K}_{\mathrm{knee}}$ by resampling evaluation queries with replacement (Appendix C).
Although $\widehat{K}_{\mathrm{knee}}$ is defined via accuracy saturation rather than Definition A.1, we use it only as an empirical baseline for comparison. To support the stability interpretation, we also directly measure distributional stability under repeated demonstration resampling on a held-out subset; see Appendix L.
We summarize prediction quality using the error ratios
$R_{\mathrm{raw}}:=\frac{K^{*}}{\widehat{K}_{\mathrm{knee}}}\qquad\text{and}\qquad R_{\mathrm{cal}}:=\frac{\widehat{K}_{\mathrm{cal}}}{\widehat{K}_{\mathrm{knee}}},$
where $K^{*}$ denotes the proxy-based prompt-length estimate produced by our observable estimator (Algorithm 1), and $\widehat{K}_{\mathrm{cal}}$ applies validation-only calibration as in (12) (Section 3.5).
This estimator is intentionally simple and nonparametric; nevertheless, to guard against spurious knees induced by noise, we (i) enforce at least two grid points on each side of the breakpoint, and (ii) report bootstrap intervals and per-seed variance statistics alongside mean ratios. Full details (grid, constraints, and bootstrapping) are provided in Appendix C.
4.3 Numerical sensitivity of spectral proxies
A practical issue is that extreme eigenvalues of $\widehat{\Sigma}_{K}$ can be highly sensitive to numerical noise (e.g., float32 accumulation, finite-sample near-singularity, and eigen-solver tolerances), which may cause large fluctuations in the predicted $K^{*}$ when using $\lambda_{\min}(\widehat{\Sigma}_{K})$ directly. We therefore test the stability of different proxies under controlled perturbations of the empirical second-moment matrix.
Protocol.
For each dataset/encoder setting, we fix demonstrations, compute $\widehat{\Sigma}_{K}$ on a grid of $K$, and add random symmetric Gaussian perturbations with relative magnitude $\varepsilon$ (scaled by $s=\mathrm{tr}(\widehat{\Sigma}_{K})/d$) to obtain perturbed second-moment matrices. For each perturbed draw we recompute the proxy-based estimate under three choices: (i) $\lambda_{\min}(\widetilde{\Sigma}_{K})$, (ii) $\lambda_{\min}(\widetilde{\Sigma}_{K}+\rho I)$ with a small ridge $\rho$, and (iii) the quantile surrogate $\lambda_{q}(\widetilde{\Sigma}_{K}+\rho I)$. We report the coefficient of variation (CV) of the resulting prompt-length estimate across perturbation draws.
Table 2: Numerical noise sensitivity of predicted prompt length $K^{*}$. Lower CV indicates higher numerical robustness.
Noise level $\varepsilon$
CV($K^{*}$ via $\lambda_{\min}$)
CV($K^{*}$ via $\lambda_{\min}+\rho I$)
CV($K^{*}$ via $\lambda_{q}$)
$10^{-6}$ 0.09 0.012 0.009
$10^{-5}$ 0.21 0.024 0.017
$10^{-4}$ 0.47 0.071 0.052
$10^{-3}$ 0.98 0.19 0.13
Analysis and mitigation via ridge regularization.
Table 2 shows that the predicted prompt length $K^{*}$ can be highly sensitive to small spectral perturbations when using $\lambda_{\min}(\widehat{\Sigma}_{K})$ directly: the CV rises from $0.09$ at $\varepsilon=10^{-6}$ to $0.98$ at $\varepsilon=10^{-3}$, indicating that the estimate becomes effectively unreliable under moderate numerical noise. Adding a small ridge $\rho$ substantially improves robustness across all noise levels (CV reduced to 0.0120.0120.190.19), consistent with the fact that ridge regularization prevents near-singularity from dominating the inverse-geometry. Specifically, we use $\widehat{\Sigma}_{K,\rho}=\widehat{\Sigma}_{K}+\rho I$ with $\rho$ set proportional to $\mathrm{tr}(\widehat{\Sigma}_{K})/d$ (see Appendix C for implementation details).
Finally, the quantile surrogate $\lambda_{q}(\widehat{\Sigma}_{K,\rho})$ is the most stable in this stress test (CV 0.0090.0090.130.13), suggesting that ignoring the extreme tail of the spectrum further reduces sensitivity while preserving the monotone relationship between spectral coverage and the required $K$.
4.4 Analysis
Table 1 summarizes the comparison between the proxy-based estimate and the empirical knee $\widehat{K}_{\mathrm{knee}}$. Across standard classification tasks, the bound is conservative yet close to scale (typical $R_{\mathrm{raw}}\approx 1.3$$1.9\times$), while validation-only calibration tightens it to within $\approx 2$$10\%$ of $1\times$ in most cases. Reasoning and RAG settings exhibit larger raw ratios, consistent with higher sensitivity to prompt perturbations, but similarly tight calibrated ratios.
Tokenization and truncation choices can measurably shift $\widehat{K}_{\mathrm{knee}}$ and the ratios; we find that single-token label verbalizers and latest-first truncation reduce knees and improve agreement (Appendix I).
Finally, controlled stress tests that violate (A1)—bounded representation drift and heavy-tailed features—match the robustness trends predicted by our extensions, and show that calibration largely preserves the proxy-based ordering under distributional shift (Appendices D and F).